# Optimal Stopping in Latent Diffusion Models

### v1

@@ A.3 Proof of Proposition 3
Thus, $t\geq T-\bar{a}^{-2}\Big(\frac{3-\varepsilon}{1+\varepsilon}\frac{\hat{\sigma}_{1}^{2}}{1-\hat{\sigma}_{1}^{2}}\Big)$ implies $t\geq\hat{t}_{d}$ for every $d\in\{1,\ldots,d_{0}\}$. On the other hand, $t<T=\hat{t}_{d}$ for all $d\in\{d_{0}+1,\ldots,D\}$ since $\sigma_{d}=\hat{\sigma}_{d}=0$. From here we deduce the desired result applying Proposition 2.
In this second part, we study under the event where $|\sigma^{2}-\hat{\sigma}_{d}^{2}|\leq\sigma^{2}$ for every $d\in\{1,\ldots,d_{0}\}$, which holds with probability $1-2d_{0}e^{-n/8}$ by Proposition 9. To prove the desired result, we first show that the minimum of the distance $d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is attained at $t=T-\hat{\delta}_{d_{0}}$, as per its definition. We consider two cases depending on whether condition (12) is satisfied. First, if condition (12) holds, the proof of Proposition 1 establishes that $a_{T-\hat{\delta}_{d_{0}}}^{2}$ is the unique zero of the derivative $\frac{d}{da_{t}^{2}}d_{F}^{2}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$. This confirms that $T-\hat{\delta}_{d_{0}}$ is the unique minimizer of the distance. Conversely, if condition (12) is not satisfied, then $\hat{\delta}_{d_{0}}=0$. In this scenario, the squared distance $d_{F}^{2}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is a non-increasing function of $t$ and thus attains its minimum at the endpoint $t=T$. This result is consistent, as $t=T=T-\hat{\delta}_{d_{0}}$.
We remark by Proposition 2 that, since $\hat{t}_{d}=T$ for every $d\in\{d_{0}+1,\ldots,D\}$, for every $t\in[0,T]$,
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{T-\hat{\delta}_{d_{0}}},\overrightarrow{X_{0}})\leq d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})\leq d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
Observe that $\hat{t}_{1}=\max_{d\in\{1,\dots,d_{0}\}}\hat{t}_{d}$, which is in the same order of $\hat{\sigma}_{d}$. This is due to the fact that $x\mapsto\frac{a-x}{1-x}$ is non-increasing if $a<1$. Then from the proof of Proposition 2 we deduce that, for $t\geq\hat{t}_{1}$ and $d\in\{1,\ldots,d_{0}\}$, that
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{T-\hat{\delta}_{d_{0}}},\overrightarrow{X_{0}})\leq d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})<d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
If $\hat{t}_{1}=0$, the proof is finished. Note that this is the case if $\sigma^{2}\geq 1/4$, since $\frac{4\sigma^{2}-\hat{\sigma}_{1}^{2}}{(1-\hat{\sigma}_{1}^{2})_{+}}\geq 1$. We study from now the case where $\hat{t}_{1}>0$ and $\sigma^{2}\leq 1/4$, with $t\leq\hat{t}_{1}$ and $d\in\{1,\ldots,d_{0}\}$. We do this by showing for every dimension $d\in\{1,\ldots,d_{0}\}$, $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is non-increasing on $0,\hat{t}_{1}]$. This implies
$d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{T-\hat{\delta}_{d_{0}}},\overrightarrow{X_{0}})\leq d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{\hat{t}_{d_{0}}},\overrightarrow{X_{0}})\leq d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{\hat{t}_{d_{0}}},\overrightarrow{X_{0}})\leq d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}}).$
In the remainder of the proof, we show, for $d\in\{1,\ldots,d_{0}\}$, that $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is non-increasing on $0,\hat{t}_{1}]$. This is equivalent to proving that $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ is non-decreasing on $T-\hat{t}_{1},T]$. Recall that, as in the proof as Proposition 1,
$d_{F}^{2}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})=\sum_{d^{\prime}=d+1}^{d_{0}}\sigma^{2}+\sum_{d^{\prime}=1}^{d}\big(\sqrt{a_{t}^{2}+(1-a_{t}^{2})\hat{\sigma}_{d^{\prime}}^{2}}-\sigma\big)^{2}.$
@@
What we want to show is equivalent to $f$ being non-decreasing on $a_{T-\hat{t}_{1}}^{2},a_{T}^{2}]$. Since $f_{d}$ is convex as proven in Proposition 1, it is sufficient to show that $f^{\prime}$ is positive at $a_{T-\hat{t}_{1}}^{2}$. All in all, since the derivative of $f_{d}$ is
@@
The result above is twofold. First, we get that $d_{F}(P_{d}^{\top}P_{d}\overleftarrow{\hat{X}}_{T-t},\overrightarrow{X_{0}})$ is increasing on the interval of interest. This also interestingly shows that the minimum of the Frobenius distance $t\mapsto d_{F}(P_{d_{0}}^{\top}P_{d_{0}}\overleftarrow{\hat{X}}_{t},\overrightarrow{X_{0}})$ is reached after $\hat{t}_{1}$. Since by definition the minium is reached at $T-\hat{\delta}_{d_{0}}$, we get that $T-\hat{\delta}_{d_{0}}\geq\hat{t}$.
@@
A.4 Proof of Proposition 4
@@ A.5 Proof of Proposition 5
A.6 Derivation of special cases of Proposition 5
First, consider the scenario where the learning capacity is unconstrained, effectively setting $C=\infty$, while the data covariance matrix is nonsingular. In this case, the condition on $d_{1}$ becomes $0\leq\hat{\sigma}_{d}^{2}$, which is trivially satisfied for all $d\in\{1,\ldots,D\}$, implying $d_{1}=D$. The condition for $d_{2}$ becomes $0>4\sigma_{d}^{2}$, which holds for none of $d$, thus implying $d_{2}=D$. Therefore, when $C=\infty$, Proposition 5 entails that $d_{\min}=D$. This result is somewhat expected: if the score function is learned perfectly, the diffusion process can be reversed in the full ambient space, enabling sampling from the target distribution without any need for dimensionality reduction.
Second, consider the scenario addressed in Proposition 3 where the true data distribution lies within a $d_{0}$-dimensional linear subspace, i.e., $\sigma_{d_{0}+1}=\cdots=\sigma_{D}=0$ and $\sigma_{1}=\cdots=\sigma_{d_{0}}=\sigma$. Assume that $C$ is sufficiently large to ensure that $1/C\leq\min(\sigma^{2},\min_{d^{\prime}\in\{1,\ldots,d_{0}\}}\hat{\sigma}_{d^{\prime}}^{2})$. Therefore, for $d\leq d_{0}$, one has $\frac{1}{C}\leq\hat{\sigma}_{d}^{2}$ (which is not satisfied anymore for $d$ beyond $d_{0}$), leading to $d_{1}=d_{0}$. On the other hand, for $d>d_{0}$ we have $\frac{1}{2C-1}\geq 0=4\sigma_{d}$. Hence $d_{0}=d_{1}\leq d_{2}\leq d_{0}$, which implies $d_{2}=d_{0}$. Thus, Proposition 5 predicts $d_{\min}=d_{0}$. This suggests that the projection onto the subspace in which the data distribution lies is the optimal sampling strategy, which is in line with the recommendation of Proposition 3.
Proof of Corollary 1.
@@ A.6 Derivation of special cases of Proposition 5
A.7 Proof of Proposition 6
The proof is concluded by using the same argument as in the proof of Proposition 2.
A.8 Proof of Proposition 7
In the case of interest for us, we get
We now argue that $\textnormal{tr}((\hat{\Sigma}_{d}^{1/2}(t)\Sigma\hat{\Sigma}_{d}^{1/2}(t))^{1/2})$ is approximately $\sum_{d^{\prime}=1}^{d}\hat{\sigma}_{d^{\prime}}\sqrt{a_{T-t}^{2}+b_{T-t}^{2}\hat{\sigma}_{d^{\prime}}^{2}}$. Observe that the two quantities are equal when $\Sigma$ and $\hat{\Sigma}$ commute, which was the case in the previous sections where we assumed that both matrices were diagonal. By Proposition 10, with probability $1-2e^{-u}$, we have $\Sigma\preceq\frac{1}{1-\varepsilon_{u}}\hat{\Sigma}$, where $\preceq$ denotes the Loewner order (see, for instance, Horn & Johnson, 2012, Definition 7.7.1). Hence,
by Lemma 2 (i). Since square root is a matrix monotonic function (see Lemma 2 (ii)), we derive that
By combining the results, we obtain
$\leq 2S(\Sigma)\varepsilon_{u}.$
Hence, for $d\in\{2,\ldots,D\}$,
We show in the following that if $t\geq\hat{T}_{d}(u)$, then
Observe that,
[…]
By squaring both sides and rearranging the terms, we deduce that
[…]
and we conclude the proof.
Remark.
We can again in this case consider the same scenario as in Proposition 3 where the eigenvalues of the covariance matrix are equal. This can be an interesting direction for future work, as to generalize the previous results to this more general setup.
Appendix B Bounds on Gaussian estimation
@@
Appendix C Useful Lemma
@@
Appendix D Experiment details
D.1 Natural image experiment
@@
We use the dataset CelebA and CelebA-HQ (Liu et al., 2015). We use a U-Net model (Ronneberger et al., 2015) and an Adam optimizer (Kingma, 2014). The diffusion model uses rectified flow noise schedule (Liu et al., 2022). The code was implemented in JAX (Bradbury et al., 2018).
Training of AE.
We train an VQ-VAE using the VQ-GAN loss (Esser et al., 2021) for $1.95$ million step on $20$ TPUv2. The VQ-VAE encodes the images to a latent space of shape $64\times 64\times 3$ and reaches an 22k-rFID score of $2.44$. Other hyperparameters for training is summarized in Table 1.
@@
Training of LDM.
@@
Noise schedule Rectified Flow
@@
Noise schedule Rectified Flow
@@
Table 3: Hyperparameters for training diffusion model on CelebA.
Training pixel diffusion model on CelebA.
We train a diffusion model on CelebA. We train for $1$ million steps on $12$ TPUv2. We summarize the hyperparameters in Table 3.
@@ Results.
Synthetic Gaussian data.

### v2

@@ Appendix B Proofs of results
Thus, $t\geq T-\bar{a}^{-2}\Big(\frac{3-\varepsilon}{1+\varepsilon}\frac{\hat{\sigma}_{1}^{2}}{1-\hat{\sigma}_{1}^{2}}\Big)$ implies $t\geq\hat{t}_{d}$ for every $d\in\{1,\ldots,d_{0}\}$. On the other hand, $t<T=\hat{t}_{d}$ for all $d\in\{d_{0}+1,\ldots,D\}$ since $\sigma_{d}=\hat{\sigma}_{d}=0$. From here we deduce the desired result applying Proposition 4.2.
[…]
We remark by Proposition 4.2 that, since $\hat{t}_{d}=T$ for every $d\in\{d_{0}+1,\ldots,D\}$, for every $t\in[0,T]$,
[…]
Observe that $\hat{t}_{1}=\max_{d\in\{1,\dots,d_{0}\}}\hat{t}_{d}$, which is in the same order of $\hat{\sigma}_{d}$. This is due to the fact that $x\mapsto\frac{a-x}{1-x}$ is non-increasing if $a<1$. Then from the proof of Proposition 4.2 we deduce that, for $t\geq\hat{t}_{1}$ and $d\in\{1,\ldots,d_{0}\}$, that
[…]
$\Delta_{d,T-t}(\hat{\Sigma})=\sum_{d^{\prime}=d+1}^{d_{0}}\sigma^{2}+\sum_{d^{\prime}=1}^{d}\big(\sqrt{a_{t}^{2}+(1-a_{t}^{2})\hat{\sigma}_{d^{\prime}}^{2}}-\sigma\big)^{2}.$
@@
What we want to show is equivalent to $f$ being non-decreasing on $a_{T-\hat{t}_{1}}^{2},a_{T}^{2}]$. Since $f_{d}$ is convex as proven in Proposition 4.1, it is sufficient to show that $f^{\prime}$ is positive at $a_{T-\hat{t}_{1}}^{2}$. All in all, since the derivative of $f_{d}$ is
@@
[…]
@@
B.4 Proof of Proposition 4.4
The proof is concluded by using the same argument as in the proof of Proposition 4.2.
B.5 Proof of Proposition 4.5
In the case of interest for us, we get
We now argue that $\textnormal{tr}((\hat{\Sigma}_{d}^{1/2}(t)\Sigma\hat{\Sigma}_{d}^{1/2}(t))^{1/2})$ is approximately $\sum_{d^{\prime}=1}^{d}\hat{\sigma}_{d^{\prime}}\sqrt{a_{T-t}^{2}+b_{T-t}^{2}\hat{\sigma}_{d^{\prime}}^{2}}$. Observe that the two quantities are equal when $\Sigma$ and $\hat{\Sigma}$ commute, which was the case in the previous sections where we assumed that both matrices were diagonal. By Proposition C.2, with probability $1-2e^{-u}$, we have $\Sigma\preceq\frac{1}{1-\varepsilon_{u}}\hat{\Sigma}$, where $\preceq$ denotes the Loewner order (see, for instance, Horn & Johnson, 2012, Definition 7.7.1). Hence,
by Lemma D.1 (i). Since square root is a matrix monotonic function (see Lemma D.1 (ii)), we derive that
By combining the results, we obtain
$\leq 2S(\Sigma)\varepsilon_{u}.$
Hence, for $d\in\{2,\ldots,D\}$,
We show in the following that if $t\geq\hat{T}_{d}(u)$, then
Observe that,
[…]
By squaring both sides and rearranging the terms, we deduce that
[…]
and we conclude the proof.
Remark.
We can again in this case consider the same scenario as in Proposition 4.3 where the eigenvalues of the covariance matrix are equal. This can be an interesting direction for future work, as to generalize the previous results to this more general setup.
B.6 Proof of Proposition 5.1
@@
B.8 Derivation of special cases of Proposition 5.2
First, consider the scenario where the learning capacity is unconstrained, effectively setting $C=\infty$, while the data covariance matrix is nonsingular. In this case, the condition on $d_{1}$ becomes $0\leq\hat{\sigma}_{d}^{2}$, which is trivially satisfied for all $d\in\{1,\ldots,D\}$, implying $d_{1}=D$. The condition for $d_{2}$ becomes $0>4\sigma_{d}^{2}$, which holds for none of $d$, thus implying $d_{2}=D$. Therefore, when $C=\infty$, Proposition 5.2 entails that $d_{\min}=D$. This result is somewhat expected: if the score function is learned perfectly, the diffusion process can be reversed in the full ambient space, enabling sampling from the target distribution without any need for dimensionality reduction.
Second, consider the scenario addressed in Proposition 4.3 where the true data distribution lies within a $d_{0}$-dimensional linear subspace, i.e., $\sigma_{d_{0}+1}=\cdots=\sigma_{D}=0$ and $\sigma_{1}=\cdots=\sigma_{d_{0}}=\sigma$. Assume that $C$ is sufficiently large to ensure that $1/C\leq\min(\sigma^{2},\min_{d^{\prime}\in\{1,\ldots,d_{0}\}}\hat{\sigma}_{d^{\prime}}^{2})$. Therefore, for $d\leq d_{0}$, one has $\frac{1}{C}\leq\hat{\sigma}_{d}^{2}$ (which is not satisfied anymore for $d$ beyond $d_{0}$), leading to $d_{1}=d_{0}$. On the other hand, for $d>d_{0}$ we have $\frac{1}{2C-1}\geq 0=4\sigma_{d}$. Hence $d_{0}=d_{1}\leq d_{2}\leq d_{0}$, which implies $d_{2}=d_{0}$. Thus, Proposition 5.2 predicts $d_{\min}=d_{0}$. This suggests that the projection onto the subspace in which the data distribution lies is the optimal sampling strategy, which is in line with the recommendation of Proposition 4.3.
Proof of Corollary 5.3.
@@
Appendix C Bounds on Gaussian estimation
@@
Appendix D Useful Lemma
@@
Appendix E Experiment details
E.1 Synthetic data experiment
Synthetic Gaussian data.
Synthetic Gaussian data with exponentially decaying eigenvalues.
In the experiment of Figure 5, we generate data using Gaussian distribution with covariance matrices equal to $\textnormal{diag}(1,1/4,1/4^{2},\ldots,1/4^{9})$. We then train a linear model with score matching using $10$ thousand samples, and we clip the parameters of the model. We generate new sample using the Ornstein-Uhlenbeck process with $T=2$ and $1000$ discretization steps.
Figure 5: Fréchet distance of the sample generated by a diffusion model, where the data follows a $10$-dimensional Gaussian distribution with exponentially decaying eigenvalues, while the score is a linear function trained with sample projected onto a lower dimension. Corollary 5.3 correctly predicts the optimal projection dimension, here $d_{\min}\in\{5,6\}$.
E.2 Natural image experiment
@@
We use the dataset CelebA and CelebA-HQ (Liu et al., 2015) and ImageNet-256 (Deng et al., 2009). We use a U-Net model (Ronneberger et al., 2015) and an Adam optimizer (Kingma, 2014). The diffusion model uses variance preserving noise schedule (Song et al., 2020). The code was implemented in JAX (Bradbury et al., 2018).
Training of AE on CelebA-HQ.
We train a VQ-VAE using the VQ-GAN loss (Esser et al., 2021) for $1.95$ million step on $20$ TPUv2. The VQ-VAE encodes the images to a latent space of shape $64\times 64\times 3$ and reaches a 22k-rFID score of $2.44$. Other hyperparameters for training is summarized in Table 1.
@@
Training of AE on ImageNet-256.
We train a VQ-VAE using the VQ-GAN loss (Esser et al., 2021) for $715$ thousand (resp. $1.81$ million) steps on $20$ TPUv2 for a latent shape of $32\times 32\times 4$ (resp. $64\times 64\times 3$). The VQ-VAE achieves an rFID score of $3.6$ (resp. $4.2$). Other hyperparameters for training is summarized in Table 2.
Name Value
Coefficient of the adversarial loss $0.1$
Coefficient of the generator loss $100$
Coefficient of the LPIPS loss $1.0$
Coefficient of the discriminator loss $0.01$
Number of embeddings of the vector quantizer $8192$
Optimizer Adam with standard hyperparameters
EMA decay $0.999$
Learning rate $10^{-5}$
Batch size $16$
Table 2: Hyperparameters for training VQ-VAE on ImageNet-256.
Training of LDM on CelebA-HQ.
@@
Noise schedule Variance Preserving
@@
Noise schedule Variance Preserving
@@
Table 4: Hyperparameters for training diffusion model on CelebA and ImageNet-64.
Training of LDM on ImageNet-256.
We train two LDMs on the images encoded by the AEs we described above. We train for $1$ million steps on $8$ TPUv6 for both LDMs. We summarize the hyperparameters used in Table 5.
Name Value
Noise schedule Variance Preserving
Number of sampling steps $250$
Optimizer Adam with standard hyperparameters
EMA decay $0.999$
Learning rate $10^{-4}$
Batch size $1024$
Table 5: Hyperparameters for training LDM on encoded images of ImageNet-256.
Training pixel diffusion model on CelebA and ImageNet-64.
We train diffusion models directly on CelebA and ImageNet-64. We train both models for $1$ million steps on $12$ TPUv2. We summarize the hyperparameters in Table 4.
@@ Results.
Linear AE and U-Net LDM on MNIST (Deng, 2012).
We train linear AE on MNIST with dimension $64$, $256$, and $400$ for 55k steps each using $4$ TPUv2. Then we train U-Net diffusion models paired with each AE for 1010k steps each using $4$ TPUv2. We summarize other hyperparameters in Table 6 and 7.
Name Value
Optimizer Adam with standard hyperparameters
Learning rate $0.003$
Batch size $256$
Table 6: Hyperparameters for training Linear AE on MNIST.
Name Value
Noise schedule Variance Preserving
Number of sampling steps $250$
Optimizer Adam with standard hyperparameters
EMA decay $0.999$
Learning rate $10^{-4}$
Batch size $256$
Table 7: Hyperparameters for training LDM on encoded images of MNIST.
Figure 6: Zoom in of Figure 3.
Figure 7: We measure the image quality of different LDMs on ImageNet-256 by MMD and Sliced Wasserstein distance.
Figure 8: We compare qualitative results of image generated with LDM and pixel diffusion. The pixel diffusion images are generated by directly adding Gaussian noise to simulate the real pixel diffusion process.
Figure 9: We train linear autoencoders on MNIST, then corresponding U-Net LDMs. The experiment shows a behavior similar to Proposition 4.2 and Figure 2: for each LDM there exists a time interval such that the LDM is optimal, and for later diffusion times, a larger latent dimension is better.
Figure 10: We train AE on CelebA-HQ and a corresponding LDM. In addition, we compare it to the FID curve of pixel diffusion model trained on CelebA. The experiment shows monotonicity in FID curve of pixel diffusion, while the FID curve of LDM achieves a minimum before the final diffusion time.